# The Properties of Little Red Dot Galaxies in the ASTRID Simulation

### v1

@@ Abstract
We present simulated counterparts of the “Little Red Dot” (LRD) galaxies observed with JWST, using the large cosmological hydrodynamic simulation, ASTRID. We create mock observations of the galaxies ($5\leq z\leq 8$) in ASTRID, and find seventeen which fit the color and size criteria of LRDs. These LRDs are galaxies with high stellar masses ($\rm log(M_{*}/\,M_{\odot})\geq 9.7$), and massive black holes ($\rm log(M_{BH}/\,M_{\odot})\geq 6.8$). The host galaxies are dense, with stellar half mass radii ($\rm 325\,pc\leq r_{{\rm half},*}\leq 620\,pc$), and dust attenuation in the F [^4] W band above 1.25. Their star formation has been recently quenched. They host relatively bright AGN that are dust-obscured and contribute significantly to the rest-frame optical red slope and have relatively low luminosity in the rest-frame ultraviolet, where the host galaxy’s stars are more dominant. These LRDs are in an evolutionary phase of miniquenching that is the result of AGN feedback from their massive black holes. The LRDs in ASTRID are bright with F [^4] W magnitudes of $23.5-25.5$. The less massive and fainter galaxies in ASTRID lack the dust concentration necessary to produce the red slope of an LRD, though this could be an effect of limited resolution. Most of the highest Eddington black holes are not LRDs due to their host galaxies having typical dust levels and relatively high star formation rates accompanying their highly accreting black holes, resulting in their spectra being too flat.
@@ 1 Introduction
One of the most remarkable and unexpected discoveries from early James Webb Space Telescope (JWST) observations has been spatially small, very red objects at high redshift ($z\geq 4$) which have been dubbed “little red dots” (LRDs; Matthee et al., 2023; Labbé et al., 2023; Kocevski et al., 2023, 2024; Harikane et al., 2023). These objects have compact or point-like morphologies (Baggen et al., 2023; Guia et al., 2024), display a red color in the observed wavelength bands between $2.0\mu m$ and $5.0\mu m$, and a flat or blue color in the $0.5\mu m$ to $2.0\mu m$ wavelength range. LRDs appear relatively common in the early universe ($z\geq 5$) with a number density of $\sim 10^{-5}\,\rm Mpc^{-3}Mag^{-1}$, implying they are not extremely rare or unique objects (Kokorev et al., 2024; Pérez-González et al., 2024; Labbé et al., 2023; Kocevski et al., 2024; Taylor et al., 2024). In this work we make mock observations of the galaxies in the cosmological simulation Astrid (Ni et al., 2022; Bird et al., 2022; Ni et al., 2024) to identify and analyze the LRD galaxy population in this simulated environment.
In addition to their compact morphology, and photometric colors, many of these LRDs have other notable qualities. These include broad Balmer line emission (Kocevski et al., 2024; Kokorev et al., 2024; Matthee et al., 2024; Greene et al., 2024), non-detection or flat spectra in X-ray and infrared (Akins et al., 2024; Ananna et al., 2024; Yue et al., 2024b; Maiolino et al., 2024a), and significant Balmer breaks (Setton et al., 2024; Rusakov et al., 2025). This combination of features has proven difficult to model, as the standard models for both stellar and AGN emission are unable to reproduce some portion of the spectrum (Durodola et al., 2025).
@@
The broad Balmer emission, combined with their red color, would support an LRD model containing a bright, but strongly dust-obscured, AGN. This model can have issues with the flat emission in the mid-IR wavelength range, and the non-detection of most LRDs in X-ray (Pacucci & Narayan, 2024), as most AGN would have a red mid-IR spectrum, and detectable X-ray emission. Additionally, many attempts to fit these objects as dominated by an AGN project the presence of black holes which are more massive than would be expected for their host galaxy and redshift, or black holes that are accreting well above the Eddington limit (Kocevski et al., 2023; Pacucci et al., 2023; Durodola et al., 2025; Taylor et al., 2024).
@@
It is also possible to model the emission of LRDs as very compact galaxies whose emission is dominated by stellar light. The stellar populations of these models fall into two broad categories: either recent significant starbursts resulting in a younger stellar population in a very dusty galaxy, or earlier starbursts producing an older stellar population with less dust and a significant Balmer break (Pérez-González et al., 2024; Williams et al., 2024; Baggen et al., 2023; Ma et al., 2024). These models are generally able to reproduce the visible spectra of LRDs, as well as the X-ray and infrared properties. Depending on the exact stellar population, dust properties, and stellar density used, some models are also able to reproduce the Balmer break, and broad emission line width of observed LRDs (Baggen et al., 2024). Early fits of stellar-dominated LRDs produced stellar masses $\gtrsim 10^{10}\,M_{\odot}$ (Akins et al., 2023; Labbé et al., 2023), placing them in the higher mass ranges for their redshift. Given their point-like appearance, they would have extremely high stellar densities, which could potentially produce the broad line widths that are seen in LRDs (Boylan-Kolchin, 2023; Akins et al., 2024; Wang et al., 2024; Endsley et al., 2023; Chworowsky et al., 2024; Baggen et al., 2024; Guia et al., 2024).
@@
Determining whether the majority of observed LRDs fall into one of these two categories, or somewhere in between, has so far proven difficult, as both the AGN and stellar proposals are able to explain some characteristics of LRDs but not yet all. Additionally, there are many important aspects of our modeling of these galaxies that are uncertain. The two most significant are modeling the AGN spectra and the galactic dust absorption, both of which play very significant roles in determining how we model or fit the observed color and brightness of a galaxy. This has become even more apparent recently as more attention has been paid to the Balmer emission and absorption features of most LRDs (D’Eugenio et al., 2025; de Graaff et al., 2025; Naidu et al., 2025; Inayoshi & Ichikawa, 2024), and the dust properties that would be necessary to produce the "V shape" visible spectrum (Setton et al., 2025). These findings suggest that an emission source with minimal dust obscuration and a very significant Balmer break, such as a gas enshrouded AGN or a compact evolved stellar population, may be the best candidate for the nature of LRDs.
@@
In this paper, we produce mock observations of galaxies in the Astrid simulation between redshifts 5 and 8 (i.e., the bulk of the redshift distribution of the LRDs, as shown in Kocevski et al. (2024)). We include the emission from both stars and AGN in order to provide a simulation comparison to the current LRD observations. In section 2, we describe the process we use to create our mock observations. In section 3, we describe the population that we produce with this mock observation pipeline, and apply LRD criteria to that population to search for LRDs. In section 4, we examine the properties of the LRDs found in Astrid, how they compare to the overall galaxy population, and possible changes that would produce different LRD populations. Finally, in section 6, we summarize the findings of our work, and discuss future work to enhance our understanding of LRD galaxies.
@@ 2 Methods
In this work, we make mock NIRCAM-like observations of galaxies in the Astrid simulation in order to identify galaxies that fit the criteria of Little Red Dots wich are described in detail in section 3. We build on the imaging process detailed in LaChance et al. (2024), adding light from AGNs, and the re-emission and dust occlusion of stellar birth clouds.
@@ 2.1 The Astrid Simulation
For this work we used the Astrid (Ni et al., 2022; Bird et al., 2022; Ni et al., 2024) simulation. Astrid  has an initial particle load of $2\times 5500^{3}$ particles in a cosmological box with side-length $250\,{\rm Mpc}h^{-1}$. This results in a mass resolution that is similar to the Illustris TNG100 and EAGLE simulations, but with a volume larger than the lower resolution Illustris TNG300 simulation (Illustris TNG: Springel et al. 2018; Nelson et al. 2018; Marinacci et al. 2018; Naiman et al. 2018; EAGLE: Schaye et al. 2015). Astrid uses values for the cosmological parameters based on the Planck survey (Planck Collaboration et al., 2020). ($\Omega_{0}=0.03089$, $\Omega_{\Lambda}=0.6911$, $\Omega_{b}=0.00486$, $\sigma_{8}=0.82$, $h=0.6774$, $n_{s}=0.9667$). Astrid has recently completed its run by reaching $z=0$.
@@
Astrid uses the star formation model described in Feng et al. (2016), developed for the BlueTides simulation, and based on the model in Springel & Hernquist (2003). Gas particles are half of the initial particle load, and can produce star particles with 1/4 of the initial mass of the gas particle. The criteria for star particle formation, and much of the gas cooling model follows the prescription in Krumholz & Gnedin (2011) and Katz et al. (1996), respectively. The full details of the handling of gas and star particles within Astrid can be found in Bird et al. (2022).
@@
The black holes in Astrid are seeded within groups of particles identified using the Friends-of-Friends algorithm (Davis et al., 1985). Once created, they are treated as particles which are subject to gravity, and several forces specific to black holes, including dynamical friction and gas drag Chen et al. (2022a). This facilitates more physically accurate trajectories and mergers than the more common approach of anchoring the seeded black hole to the galaxy’s minimum potential (Chen et al., 2022b). Black hole accretion is calculated with a Bondi-Hoyle-Lyttleton style model (Di Matteo et al., 2005; Bondi, 1952; Bondi & Hoyle, 1944; Hoyle & Lyttleton, 1939). Our model allows for accretion to exceed the Eddington accretion rate, with a hard limit at twice the Eddington rate. The thermal feedback that results from AGN accretion is also modeled by imparting energy into the gas particles near the AGN at a rate proportional to the black hole’s accretion. The full description of the handling of black holes in the Astrid simulation is detailed in Ni et al. (2022).
@@
In order to identify and analyze galaxies within the simulation, we must identify substructures within the Friends-of-Friends groups. The subhalo identification algorithm SUBFIND (Springel et al., 2001) is then run to identify galaxy-like substructures in these friends-of-friends groups. The properties and population statistics of these subhalos can be found in Bird et al. (2022), which showed that subhalo statistics are in good agreement with observational data.
@@
For this work, we focused on the snapshots at redshifts z=5, 6, 7, and 8 as that covers the majority of the redshift range where LRDs have been seen in observations (Kocevski et al., 2024). We analyze all subhalos in these snapshots with a stellar mass above $10^{8}\,M_{\odot}$ as this is near the lowest stellar mass fit from observations. Unlike LaChance et al. (2024), we do not perform any cuts for observation quality, as we are not analyzing the morphologies of the entire population, and are instead focused on the photometry of the galaxies.
@@ 2.2 Galaxy imaging process
We use the imaging pipeline described in LaChance et al. (2024) as the starting point for our imaging process in order to produce mock observations like those seen in figure 1. This pipeline uses the star and gas particles present in subhalos found using SUBFIND to calculate the intrinsic stellar emission, and the dust occlusion for each star particle. We use the Binary Population and Spectral Population Synthesis model (BPASS version 2.2.1; Stanway & Eldridge, 2018) to calculate the intrinsic stellar emission for each star based on its age, mass, and metallicity. The dust occlusion due to the interstellar medium (ISM) for each star at a given wavelength is calculated using the following equation
@@
where $\tau_{\rm ISM}(\lambda)$ is the optical depth at a given wavelength, $\kappa_{\rm ISM}$ is a tuning parameter, which we assign a value of $10^{4.1}$ based on the normalization in Bird et al. (2022), $\Sigma(x,y,z)$ is the metal surface density for that star along the line of sight, and $\gamma$ is a tuning parameter for the slope of the dust model, which we assign a value of $-1.0$ which falls between the slopes of the Small Magellanic Cloud model (Pei, 1992) and the ‘starburst’ model (Calzetti et al., 2000). For further details, see Wilkins et al. (2017); LaChance et al. (2024).
@@
We make two additions to this imaging process in order to emulate emission contributions that are vital to LRDs but were not included in the pipeline used in LaChance et al. (2024). First, we add the light from galaxy AGNs using spectra produced with CLOUDY. Second, we include the re-emission, and dust occlusion from stellar birth clouds to our stellar dust occlusion model.
@@ 2.2.1 AGN SED calculation
We use CLOUDY to create our intrinsic AGN SEDs. We use the CLOUDY default AGN input SED, with the default parameters of $-1.4$ for $\alpha_{\rm ox}$, $-0.5$ for $\alpha_{\rm UV}$, and $-1.0$ for $\alpha_{X}$. We produce 14 different SEDs with temperatures ranging from $10^{4.6}$ K to $10^{6.0}$ K in order to capture the range of temperatures present in the Astrid simulation in our redshift range. We produce these SEDs at a set bolometric luminosity of $10^{43.5}\rm erg/s$, and with the following cloudy parameters: inner hydrogen density $=10^{9}$ $\rm N/cm^{-3}$, inner radius $=5\times 10^{16}$ cm, total hydrogen column density $=10^{23}$ $\rm N/cm^{2}$, and metallic abundance of 1/10th the old solar abundance preset.
@@
Figure 3: Color-magnitude, and color-color plots showing the main criteria for LRD identification. The red circles indicate galaxies that fit both criteria, while the green squares are galaxies that fit the F277W-F [^4] W criteria, but not the F150W-F200W criteria. The contours contain $1\sigma$ and $2\sigma$ of galaxies in their subpopulation.
We use equation 1 for the dust occlusion of our black holes as well. That said, we change the value of $\gamma$ from $-1.0$ to $-2.0$, producing a dust extinction curve that is more sensitive to wavelength. There is evidence of differences in AGN and stellar dust attenuation laws in LRDs (Brooks et al., 2024; Killi et al., 2024), and our specific choice of $\gamma=-2.0$ does align with the dust model used in Ma et al. (2024). Additionally, we calculate the metal surface density for the black hole on its precise line of sight, using the SPH kernel of each gas particle to calculate its contribution. Specifically, we follow the method described in Ni et al. (2022). We compared this method to the method used for the star particles as described in LaChance et al. (2024) and found minimal difference between the two, with the SPH kernel method being more accurate in cases with very dense, low smoothing length gas particles.

### v2

@@ Abstract
We present simulated counterparts of the “Little Red Dot” (LRD) galaxies observed with JWST, using the large cosmological hydrodynamic simulation, ASTRID. We create mock observations of the galaxies ($5\leq z\leq 8$) in ASTRID, and find seventeen which fit the color and size criteria of LRDs. These LRDs are galaxies with high stellar masses ($\rm log(M_{*}/\,M_{\odot})\geq 9.7$), and massive black holes ($\rm log(M_{BH}/\,M_{\odot})\geq 6.8$). The host galaxies are dense, with stellar half mass radii ($\rm 325\,pc\leq r_{{\rm half},*}\leq 620\,pc$), and dust attenuation in the F [^4] W band above 1.25. Their star formation has been recently quenched. They host relatively bright AGN that are dust-obscured and contribute significantly to the rest-frame optical red slope and have relatively low luminosity in the rest-frame ultraviolet, where the host galaxy’s stars are more dominant. These LRDs are in an evolutionary phase of miniquenching that is the result of AGN feedback from their massive black holes. The LRDs in ASTRID are bright with F [^4] W magnitudes of 23.525.523.5-25.5. The less massive and fainter galaxies in ASTRID lack the dust concentration necessary to produce the red slope of an LRD, though this could be an effect of limited resolution. Most of the highest Eddington black holes are not LRDs due to insufficient dust attenuation from their host galaxies, which are also experiencing relatively high star formation rates. This results in their spectra being too flat, despite their highly accreting black holes.
@@ 1 Introduction
One of the most remarkable and unexpected discoveries from early James Webb Space Telescope (JWST) observations has been spatially small, very red objects at high redshift ($z\geq 4$) which have been dubbed “little red dots” (LRDs; Matthee_2023; Labbe_2023; Kocevski_2023; Kocevski_2024; Harikane_2023). These objects have compact or point-like morphologies (Baggen_2023; Guia_2024), display a red color in the observed wavelength bands between $2.0\mu m$ and $5.0\mu m$, and a flat or blue color in the $0.5\mu m$ to $2.0\mu m$ wavelength range. LRDs appear relatively common in the early universe ($z\geq 5$) with a number density of $\sim 10^{-5}\,\rm Mpc^{-3}Mag^{-1}$, implying they are not extremely rare or unique objects (Kokorev_2024; Perez_Gonzalez_2024; Labbe_2023; Kocevski_2024; Taylor_2024). In this work we make mock observations of the galaxies in the cosmological simulation Astrid (astrid_BHs; astrid_galaxy_formation; Astrid2024) to identify and analyze the LRD galaxy population in this simulated environment.
In addition to their compact morphology, and photometric colors, many of these LRDs have other notable qualities. These include broad Balmer line emission (Kocevski_2024; Kokorev_2024; matthee_2024; Greene_2024), non-detection or flat spectra in X-ray and infrared (Akins_2024; Ananna_2024; Yue_2024; Maiolino_2024), and significant Balmer breaks (Setton_2024; Rusakov_2025). This combination of features has proven difficult to model, as the standard models for both stellar and AGN emission are unable to reproduce some portion of the spectrum (Durodola_2024).
@@
The broad Balmer emission, combined with their red color, would support an LRD model containing a bright, but strongly dust-obscured, AGN. This model can have issues with the flat emission in the mid-IR wavelength range, and the non-detection of most LRDs in X-ray (Pacucci_Narayan_2024), as most AGN would have a red mid-IR spectrum, and detectable X-ray emission. Additionally, many attempts to fit these objects as dominated by an AGN project the presence of black holes which are more massive than would be expected for their host galaxy and redshift, or black holes that are accreting well above the Eddington limit (Kocevski_2023; Pacucci_2023; Durodola_2024; Taylor_2024).
Recent works in this area have found that models which embed the AGN in a very dense gas cloud can also reproduce some aspects of the LRD spectrum, while avoiding many of the issues associated with other AGN models (Naidu_2025; Taylor_2025; Inayoshi_2025; Jeon_2025). These “gas-enshrouded” AGN models produce the rest-visible red color of LRDs via a very strong Balmer break. This can alleviate the non-detection of LRDs in both the X-ray and IR wavelengths. The dense enshrouding gas which creates the Balmer break, also provides a significant amount of attenuation to any X-ray emission, potentially lowering the resulting X-ray flux below current detection limits. Additionally, the lack of dust attenuation as a source of reddening reduces the amount of dust re-emission that would be expected in the IR bands, which may alleviate tensions with observational detection limits in this region as well.
@@
It is also possible to model the emission of LRDs as very compact galaxies whose emission is dominated by stellar light. The stellar populations of these models fall into two broad categories: either recent significant starbursts resulting in a younger stellar population in a very dusty galaxy, or earlier starbursts producing an older stellar population with less dust and a significant Balmer break (Perez_Gonzalez_2024; Williams_2024; Baggen_2023; Ma_2024). These models are generally able to reproduce the visible spectra of LRDs, as well as the X-ray and infrared properties. Depending on the exact stellar population, dust properties, and stellar density used, some models are also able to reproduce the Balmer break, and broad emission line width of observed LRDs (Baggen_2024). Early fits of stellar-dominated LRDs produced stellar masses $\gtrsim 10^{10}\,M_{\odot}$ (Akins_2023; Labbe_2023), placing them in the higher mass ranges for their redshift. Given their point-like appearance, they would have extremely high stellar densities, which could potentially produce the broad line widths that are seen in LRDs (Boylan-Kolchin_2023; Akins_2024; Wang_2024; Endsley_2023; Chworowsky_2024; Baggen_2024; Guia_2024).
@@
Determining whether the majority of observed LRDs fall into one of these two categories, or somewhere in between, has so far proven difficult, as both the AGN and stellar proposals are able to explain some characteristics of LRDs but not yet all. Additionally, there are many important aspects of our modeling of these galaxies that are uncertain. The two most significant are modeling the AGN spectra and the galactic dust absorption, both of which play very significant roles in determining how we model or fit the observed color and brightness of a galaxy. This has become even more apparent recently as more attention has been paid to the Balmer emission and absorption features of most LRDs (DEugenio_2025; deGraff_2025; Naidu_2025; Inayoshi_2024), and the dust properties that would be necessary to produce the "V shape" visible spectrum (Setton_2025). These findings suggest that an emission source with minimal dust obscuration and a very significant Balmer break, such as a gas enshrouded AGN or a compact evolved stellar population, may be the best candidate for the nature of LRDs.
@@
In this paper, we produce mock observations of galaxies in the Astrid simulation between redshifts 5 and 8 (i.e., the bulk of the redshift distribution of the LRDs, as shown in Kocevski_2024). We include the emission from both stars and AGN in order to provide a simulation comparison to the current LRD observations. In section 2, we describe the process we use to create our mock observations. In section 3, we describe the population that we produce with this mock observation pipeline, and apply LRD criteria to that population to search for LRDs. In section 4, we examine the properties of the LRDs found in Astrid, how they compare to the overall galaxy population, and possible changes that would produce different LRD populations. Finally, in section 6, we summarize the findings of our work, and discuss future work to enhance our understanding of LRD galaxies.
@@ 2 Methods
In this work, we make mock NIRCAM-like observations of galaxies in the Astrid simulation in order to identify galaxies that fit the criteria of LRDs which are described in detail in section 3. We build on the imaging process detailed in Lachance_24, adding light from AGNs, and the re-emission and dust occlusion of stellar birth clouds.
@@ 2.1 The Astrid Simulation
For this work we used the Astrid (astrid_BHs; astrid_galaxy_formation; Astrid2024) simulation. Astrid has an initial particle load of $2\times 5500^{3}$ particles in a cosmological box with side-length $250\,{\rm Mpc}h^{-1}$. This results in a dark matter mass resolution of ($\rm M_{DM}=9.63\times 10^{6}\,M_{\odot}$) that is similar to the Illustris TNG100($\rm M_{DM}=7.5\times 10^{6}\,M_{\odot}$) and EAGLE ($\rm M_{DM}=9.7\times 10^{6}\,M_{\odot}$) simulations, but with a volume larger than the lower resolution Illustris TNG300 simulation (Illustris TNG: Springel_2018; Nelson_2018; Marinacci_2018; Naiman_2018; EAGLE: Schaye_2015). Astrid uses values for the cosmological parameters based on the Planck survey (Planck_2020). ($\Omega_{0}=0.03089$, $\Omega_{\Lambda}=0.6911$, $\Omega_{b}=0.00486$, $\sigma_{8}=0.82$, $h=0.6774$, $n_{s}=0.9667$). Astrid has recently completed its run by reaching $z=0$.
@@
Astrid uses the star formation model described in Feng_2016, developed for the BlueTides simulation, and based on the model in Springel_2003. Gas particles are half of the initial particle load, and can produce star particles with 1/4 of the initial mass of the gas particle. The criteria for star particle formation, and much of the gas cooling model follows the prescription in Krumholz_2011; and Katz_1996, respectively. The full details of the handling of gas and star particles within Astrid can be found in astrid_galaxy_formation.
@@
The black holes in Astrid are seeded within groups of particles identified using the Friends-of-Friends algorithm (Davis_1985). Once created, they are treated as particles which are subject to gravity, and several forces specific to black holes, including dynamical friction and gas drag Chen_2022. This facilitates more physically accurate trajectories and mergers than the more common approach of anchoring the seeded black hole to the galaxy’s minimum potential (Chen_2022b). Black hole accretion is calculated with a Bondi-Hoyle-Lyttleton style model (DiMatteo_2005; Bondi_1952; Bondi_1944; Hoyle_1939). Our model allows for accretion to exceed the Eddington accretion rate, with a hard limit at twice the Eddington rate. The thermal feedback that results from AGN accretion is also modeled by imparting energy into the gas particles near the AGN at a rate proportional to the black hole’s accretion. The full description of the handling of black holes in the Astrid simulation is detailed in astrid_BHs.
@@
In order to identify and analyze galaxies within the simulation, we must identify substructures within the Friends-of-Friends groups. The subhalo identification algorithm SUBFIND (Springel_2001) is then run to identify galaxy-like substructures in these friends-of-friends groups. The properties and population statistics of these subhalos can be found in astrid_galaxy_formation, which showed that subhalo statistics are in good agreement with observational data.
@@
For this work, we focused on the snapshots at redshifts z=5, 6, 7, and 8 as that covers the majority of the redshift range where LRDs have been seen in observations (Kocevski_2024). We analyze all subhalos in these snapshots with a stellar mass above $10^{8}\,M_{\odot}$ as this is near the lowest stellar mass fit from observations. Unlike Lachance_24, we do not perform any cuts for observation quality, as we are not analyzing the morphologies of the entire population, and are instead focused on the photometry of the galaxies.
@@ 2.2 Galaxy imaging process
We use the imaging pipeline described in Lachance_24 as the starting point for our imaging process in order to produce mock observations like those seen in figure 1. This pipeline uses the star and gas particles present in subhalos found using SUBFIND to calculate the intrinsic stellar emission, and the dust occlusion for each star particle. We use the Binary Population and Spectral Population Synthesis model (BPASS version 2.2.1; Stanway2018) to calculate the intrinsic stellar emission for each star based on its age, mass, and metallicity. The dust occlusion due to the interstellar medium (ISM) for each star at a given wavelength is calculated using the following equation
@@
where $\tau_{\rm ISM}(\lambda)$ is the optical depth at a given wavelength, $\kappa_{\rm ISM}$ is a tuning parameter, which we assign a value of $10^{4.1}$ based on the normalization in astrid_galaxy_formation, $\Sigma(x,y,z)$ is the metal surface density for that star along the line of sight, and $\gamma$ is a tuning parameter for the slope of the dust model, which we assign a value of 1.0-1.0 which falls between the slopes of the Small Magellanic Cloud model (Pei_1992) and the ‘starburst’ model (Calzetti_2000). For further details, see Wilkins17; Lachance_24.
@@
We make two additions to this imaging process in order to emulate emission contributions that are vital to LRDs but were not included in the pipeline used in Lachance_24. First, we add the light from galaxy AGNs using spectra produced with CLOUDY (cloudy). Second, we include the re-emission, and dust occlusion from stellar birth clouds to our stellar dust occlusion model.
@@ 2.2.1 AGN SED calculation
We use CLOUDY to create our intrinsic AGN SEDs. We use the CLOUDY default AGN input SED, with the default parameters of 1.4-1.4 for $\alpha_{\rm ox}$, 0.5-0.5 for $\alpha_{\rm UV}$, and 1.0-1.0 for $\alpha_{X}$. We produce 14 different SEDs with temperatures ranging from $10^{4.6}$ K to $10^{6.0}$ K in order to capture the range of temperatures present in the Astrid simulation in our redshift range. We produce these SEDs at a set bolometric luminosity of $10^{43.5}\rm erg/s$, and with the following cloudy parameters: inner hydrogen density $10^{9}$ $\rm N/cm^{-3}$, inner radius $5\times 10^{16}$ cm, total hydrogen column density $10^{23}$ $\rm N/cm^{2}$, and metallic abundance of 1/10th the old solar abundance preset.
@@
This process models the emission from both the narrow and broad line regions (NLR, BLR) surrounding the AGN, with the inner radius of the gas cloud representing the inner radius of the BLR. Notably, the gas cloud in our model does not include the dynamics of the BLR, which prevents the broadening of emission lines produced in that region. That emission is still included in our model, but is concentrated at the wavelength of the emission line rather than being broadened.
@@
We incorporate as much information from the black holes in the simulation as possible when calculating their SED. This includes their mass and accretion rate as described above. Unfortunately, the properties of the surrounding gas cloud are not well defined in the simulation due to their small scale relative to the resolution of the simulation. As a result, we must make some assumptions about these properties. While we choose gas properties that are consistent with standard AGN models, there are recent works which suggest the possibility of significantly denser gas clouds, which would significantly alter the transmitted AGN spectrum (Naidu_2025; Taylor_2025; Inayoshi_2025; Jeon_2025). We perform an exploratory analysis of a similar model in section 5, and conduct an in depth analysis of a similar model in (LaChance et al. In prep.).
Figure 3: Color-magnitude, and color-color plots showing the main criteria for LRD identification. The red circles indicate galaxies that fit both criteria, while the green squares are galaxies that fit the F277W-F [^4] W criteria, but not the F150W-F200W criteria. The contours contain $1\sigma$ and $2\sigma$ of galaxies in their subpopulation.. Observational data sets from Perez_Gonzalez_2024; Barro_2023; and Labbe_2023 are included as grey stars, squares, and triangles respectively
We use equation 1 for the dust occlusion of our black holes as well. That said, we change the value of $\gamma$ from 1.0-1.0 to 2.0-2.0, producing a dust extinction curve that is more sensitive to wavelength. There is evidence of differences in AGN and stellar dust attenuation laws in LRDs (Brooks_2024; Killi_2024), and our specific choice of $\gamma=-2.0$ does align with the dust model used in Ma_2024. Additionally, we calculate the metal surface density for the black hole on its precise line of sight, using the SPH kernel of each gas particle to calculate its contribution. Specifically, we follow the method described in astrid_BHs. We compared this method to the method used for the star particles as described in Lachance_24 and found minimal difference between the two, with the SPH kernel method being more accurate in cases with very dense, low smoothing length gas particles.